# brick-splitting Torsion Pairs and Trim Lattices

### v1

@@ Abstract.
We introduce the notion of brick-splitting torsion pairs as a modern analogue and generalization of the classical notion of splitting torsion pairs. A torsion pair is called brick-splitting if any given brick is either torsion or torsion-free with respect to that torsion pair. After giving some properties of these pairs, we fully characterize them in terms of some lattice-theoretical properties, including left modularity. This leads to the notion of brick-directed algebras, which are those for which there does not exist any cycle of non-zero non-isomorphisms between bricks. This class of algebras is a novel generalization of representation-directed algebras. We show that brick-directed algebras have many interesting properties and give several characterizations of them. In particular, we prove that a brick-finite algebra is brick-directed if and only if the lattice of torsion classes is a trim lattice. We also give a characterization of brick-directed algebras in terms of their wall-and-chamber structure, as well as of a certain Newton polytope associated to them. Moreover, we introduce an explicit construction of an abundance of brick-directed algebras, both of the tame and wild representation types.
@@ 1.1. Brick-splitting torsion pairs
Theorem 1.1 (Prop. 3.12 &\&& Prop. 3.15)
@@
Motivated by this observation, we say an algebra $A$ is fully cyclic provided that for any idempotent $e$ of $A$, the condition $eA(1-e)=0$ implies $e=0$ or $e=1$. The following corollary is an interesting consequence of Proposition 1.2.
Corollary 1.3 (Cor. 3.9)
@@
As our second novel generalization of a classical notion in representation theory, we give a brick-analogue of the family of representation-directed algebras (for the definition and some remarks on representation-directed algebras, see Section 4). For an algebra $A$, we say $X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}\cdots\xrightarrow{f_{m-1}}X_{m-1}\xrightarrow{f_{m}}X_{m}$ is a brick-cycle in $\operatorname*{mod}A$ if every $X_{i}$ is a brick, for all $0\leq i\leq m$, with $X_{0}\simeq X_{m}$, and each $f_{j}:X_{j-1}\rightarrow X_{j}$ is non-zero and non-invertible, for $1\leq j\leq m$. Then, we say $A$ is a brick-directed algebra if $\operatorname*{mod}A$ contains no brick-cycles. Evidently, each representation-directed algebra is brick-directed. However, there are many brick-directed algebras that are not representation-directed (see Examples 3.5, 4.14, and Corollary 4.16).
@@
1.2. $\tau$-tilting finite algebras and trim lattices
@@
Recall that a finite lattice $(L,\leq)$ is said to be extremal if it has a maximal chain whose length is equal to both the number of join-irreducible elements and the number of meet-irreducibles. Extremal lattices were first introduced in [Ma], as a generalization of distributive lattices. Another generalization of distributive lattices is given by the left modularity property: $(L,\leq)$ is called left modular if $L$ has a maximal chain of left modular elements. Note that these two generalizations of distributive lattices are independent, that is, neither of them implies the other (see Section 2 and references therein). Following [Th], a finite lattice $(L,\leq)$ is said to be trim if it is extremal and left modular. Trim lattices were introduced as an analogue of distributive lattices, while they are not necessarily graded. In fact, $L$ is distributive if and only if it is graded and trim ([Th, Theorem 2]). As shown more recently in [TW], trim lattices admit many interesting combinatorial properties.
To put the following theorem into perspective, let us make some observations. We first recall that for any algebra $A$, the lattice $\operatorname*{tors}A$ is always semidistributive (see [DI+, Theorem 1.3] and [RST, Corollary 8.8]). Meanwhile, an arbitrary finite semidistributive lattice need not be left modular nor extremal. Moreover, there are trim lattices which are not semidistributive (see [TW, Figure 3]). Through some lattice-theoretical techniques, it has recently been shown that for semidistributive lattices, left modularity, extremality, and trimness are equivalent (see [TW] and [Mu]). However, there is no full classification of those algebras $A$ for which $\operatorname*{tors}(A)$ is a trim lattice. The following theorem gives such a complete classification result. Moreover, we use some properties of brick-splitting torsion pairs to give new realizations of the above-mentioned lattice-theoretical properties in the more algebraic framework of representation theory (see Propositions 5.1 and 5.2). Before we summarize our results in the next theorem, recall that for an extremal lattice $L$, the spine of $L$ consists of those elements that lie on a chain of maximal length.
Theorem 1.7 (Prop. 5.1 &\&& Prop. 5.2 &\&& Cor. 5.3)
@@
Let $A$ be a brick-finite algebra. If $A$ is brick-directed, then each brick in $\operatorname*{mod}A$ is uniquely determined by its dimension vector, that is, for two distinct elements $X$ and $Y$ in $\operatorname*{brick}A$, we have $\underline{\dim}X\neq\underline{\dim}Y$.
@@
Let $A$ be brick-finite, and set $M$ as the direct sum of $X\in\operatorname*{brick}A$. Then, $A$ is brick directed if and only if there exists an indivisible increasing path in $\mathrm{N}(M)$ from 00 to $[M]$.
@@ 2.1. Lattice theory
In the following, $(L,\leq)$ always denotes a complete lattice, and $1$ and 00 respectively denote the greatest element and the least element of $L$ (also known as the maximum and minimum of $L$). Whenever there is no confusion, we suppress $\leq$ in the notation and simply use $L$ instead of $(L,\leq)$. Recall that a lattice $L$ is said to be weakly atomic if for each $x<y$ in $L$, there is at least one arrow in $\operatorname*{Hasse}[x,y]$.
@@
An element $x\in L$ is called join-irreducible if, for each $y$ and $z$ in $L$ with $y,z<x$, we have $x\neq y\vee z$. By convention, 00 is not considered as a join-irreducible element. Provided $L$ is a finite lattice, $x$ is join-irreducible if and only if it covers exactly one element of $L$. Moreover, $x\in L$ is said to be completely join-irreducible if for any subset $S\subseteq\{y\in L\mid y<x\}$, we have $x\neq\bigvee_{y\in S}y$. By $\operatorname*{JI}(L)$ we denote the set of all join-irreducible elements in $L$, and $\operatorname*{JI}^{c}(L)$ denotes the subset of $\operatorname*{JI}(L)$ consisting of completely join-irreducible elements. The notion of (completely) meet-irreducible elements, and consequently the sets $\operatorname*{MI}(L)$ and $\operatorname*{MI}^{c}(L)$, are defined dually. Note that, by convention, $1$ does not belong to $\operatorname*{MI}(L)$.
@@
Since the extremal and left modular lattices give two interesting generalizations of the finite distributive lattices, it is natural to ask about the intersection of these two families. Following [Th], we say $L$ is a trim lattice if $L$ is simultaneously extremal and left modular. The family of trim lattices properly contains that of finite distributive lattices, yet they manifest some important properties similar to those of distributive lattices. In particular, $L$ is distributive if and only if $L$ is a graded trim lattice ([Th, Theorem 2]). More recently, new properties of trim lattices have been studied in [TW].
Before we recall some important results of [TW] and [Mu] that relate to our work, we note that the notions of extremality and left modularity do not imply each other. More explicitly, there exist extremal lattices that are not left modular, and there also exist left modular lattices that are not extremal (see [TW, Figure 5]). Moreover, as shown in [TW, Figure 3], there exist trim lattices that are not semidistributive, and there also exist semidistributive lattices that are not trim.
Theorem 2.1 ([TW, Theorem 1.4])
Let $(L,\leq)$ be a finite lattice. If $L$ is extremal and semidistributive, then $L$ is left modular. That being the case, $L$ is a trim lattice.
Inspired by the above theorem, in [Mu] the author conducted a thorough study of left modular semidistributive lattices in a slightly more general setting. In particular, a consequence of his main theorem, the converse of Theorem 2.1 is obtained, which we state in the following theorem.
Theorem 2.2 ([Mu, Theorem 3.2])
Let $(L,\leq)$ be a finite lattice. If $L$ is left modular and semidistributive, then $L$ is extremal. That being the case, $L$ is a trim lattice.
From the last two theorems and definition of trim lattices, it is immediate that if $(L,\leq)$ is a finite semidistributive lattice, then the following are equivalent:
@@ 2.2. Torsion theories
    $\mathcal{T}$ and $\mathcal{F}$ have only 00 in common.
@@ 2.4. Chain of bricks
We note that the map $f$ in the above theorem is almost never invertible. In fact, $f$ is bijective if and only if $A$ is local. Moreover, one immediately concludes that the length of each chain in $\operatorname*{tors}A$ is bounded by $|{\operatorname*{brick}A}|$. Also, note that $A$ is brick-infinite if and only if $\operatorname*{tors}A$ contains an infinite chain. For more details on the chain of torsion classes and their interactions with bricks, see [BCZ], [DIJ], [DI+], and [KD, Appendix].

### v2

# Brick-splitting torsion pairs and left modularity

@@ Abstract.
We introduce the notion of brick-splitting torsion pairs as a modern analogue and generalization of the classical notion of splitting torsion pairs. A torsion pair is called brick-splitting if any given brick is either torsion or torsion-free with respect to that torsion pair. After giving some properties of these pairs, we fully characterize them in terms of some lattice-theoretical properties, including left modularity. This leads to the notion of brick-directed algebras, which are those for which there does not exist any cycle of non-zero non-isomorphisms between bricks. This class of algebras is a novel generalization of representation-directed algebras. We show that brick-directed algebras have many interesting properties and give several characterizations of them. In particular, we prove that a brick-finite algebra is brick-directed if and only if the lattice of torsion classes is left modular (or equivalently, extremal). We also give a characterization of brick-directed algebras in terms of their wall-and-chamber structure, as well as of a certain Newton polytope associated to them. Moreover, we introduce an explicit construction of an abundance of brick-directed algebras, both of the tame and wild representation types.
@@ 1.1. Brick-splitting torsion pairs
Theorem 1.1 (Prop. 3.1 &\& Prop. 3.4)
@@
Motivated by this observation, we say an algebra $A$ is fully cyclic provided that for any idempotent $e$ of $A$, the condition eA(1e)=0eA(1-e)=0 implies $e=0$ or $e=1$. The following corollary is an interesting consequence of Proposition 1.2.
Corollary 1.3 (Cor. 3.16)
@@
As our second novel generalization of a classical notion in representation theory, we give a brick-analogue of the family of representation-directed algebras (for the definition and some remarks on representation-directed algebras, see Section 4). For an algebra $A$, we say $X_{0}\xrightarrow{f_{1}}X_{1}\xrightarrow{f_{2}}\cdots\xrightarrow{f_{m-1}}X_{m-1}\xrightarrow{f_{m}}X_{m}$ is a brick-cycle in $\operatorname*{mod}A$ if every $X_{i}$ is a brick, for all $0\leq i\leq m$, with $X_{0}\simeq X_{m}$, and each $f_{j}:X_{j-1}\rightarrow X_{j}$ is non-zero and non-invertible, for $1\leq j\leq m$. Then, we say $A$ is a brick-directed algebra if $\operatorname*{mod}A$ contains no brick-cycles. Evidently, each representation-directed algebra is brick-directed. However, there are many brick-directed algebras that are not representation-directed (see Examples 3.12, 4.11, and Corollary 4.13).
@@
Notice that, by considering a particular class of bricks, a similar (but weaker) notion to brick-directed algebras (algebras admitting a reverse hom-orthogonal order) was studied in [BH, Definition 5.5]. We observe that for brick-finite algebras, the aforementioned notion coincides with the notion of brick-directedness that we introduced above.
1.2. $\tau$-tilting finite algebras and left modular lattices
@@
We first recall that for any algebra $A$, the lattice $\operatorname*{tors}A$ is always semidistributive (see [DI+, Theorem 1.3] and [RST, Corollary 8.8]), a weaker version of distributive lattices. We also recall that, in lattice theory, there are some other weaker versions of distributive lattices. Firstly, $(L,\leq)$ is called left modular if $L$ has a maximal chain of left modular elements. Secondly, $(L,\leq)$ is said to be extremal if it has a maximal chain whose length is equal to both the number of join-irreducible elements and the number of meet-irreducibles [Ma]. Thirdly, a finite lattice $(L,\leq)$ is said to be trim if it is extremal and left modular [Th]. Through some lattice-theoretical techniques, it has recently been shown that for semidistributive lattices, left modularity, extremality, and trimness are equivalent (see [TW] and [Mu]). However, without semidistributivity, neither of left modularity and extremality implies the other (see Section 2 and references therein).
The following theorem gives new realizations of the above-mentioned lattice-theoretical properties in the more algebraic framework of representation theory (see Propositions 5.1 and 5.2). In particular, we give a complete classification of those algebras $A$ for which $\operatorname*{tors}A$ are left modular (equivalently, extremal). Before we summarize our results in the next theorem, recall that for an extremal lattice $L$, the spine of $L$ consists of those elements that lie on some chain of maximal length.
Theorem 1.7 (Prop. 5.1 &\& Prop. 5.2 &\& Cor. 5.3)
@@
Let $A$ be a brick-finite algebra. If $A$ is brick-directed, then each brick in $\operatorname*{mod}A$ is uniquely determined by its dimension vector, that is, for two distinct elements $X$ and $Y$ in $\operatorname*{brick}A$, we have $\operatorname{\underline{dim}}X\neq\operatorname{\underline{dim}}Y$.
@@
Let $A$ be brick-finite, and set $M$ as the direct sum of $X\in\operatorname*{brick}A$. Then, $A$ is brick directed if and only if there exists an indivisible increasing path in $\mathrm{N}(M)$ from 0 to $[M]$.
@@ 2.1. Lattice theory
In the following, $(L,\leq)$ always denotes a complete lattice, and $1$ and 0 respectively denote the greatest element and the least element of $L$ (also known as the maximum and minimum of $L$). Whenever there is no confusion, we suppress $\leq$ in the notation and simply use $L$ instead of $(L,\leq)$. Recall that a lattice $L$ is said to be weakly atomic if for each $x<y$ in $L$, there is at least one arrow in $\operatorname*{Hasse}[x,y]$.
@@
An element $x\in L$ is called join-irreducible if, for each $y$ and $z$ in $L$ with $y,z<x$, we have $x\neq y\vee z$. By convention, 0 is not considered as a join-irreducible element. Provided $L$ is a finite lattice, $x$ is join-irreducible if and only if it covers exactly one element of $L$. Moreover, $x\in L$ is said to be completely join-irreducible if for any subset $S\subseteq\{y\in L\mid y<x\}$, we have $x\neq\bigvee_{y\in S}y$. By $\operatorname*{JI}(L)$ we denote the set of all join-irreducible elements in $L$, and $\operatorname*{JI}^{c}(L)$ denotes the subset of $\operatorname*{JI}(L)$ consisting of completely join-irreducible elements. The notion of (completely) meet-irreducible elements, and consequently the sets $\operatorname*{MI}(L)$ and $\operatorname*{MI}^{c}(L)$, are defined dually. Note that, by convention, $1$ does not belong to $\operatorname*{MI}(L)$.
@@
Since the extremal and left modular lattices give two interesting generalizations of the finite distributive lattices, it is natural to ask about the intersection of these two families. Following [Th], we say $L$ is a trim lattice if $L$ is simultaneously extremal and left modular. The family of trim lattices properly contains that of finite distributive lattices, yet they manifest some important properties similar to those of distributive lattices. In particular, $L$ is distributive if and only if $L$ is a graded trim lattice ([Th, Theorem 2]). As shown more recently in [TW], trim lattices admit many interesting combinatorial properties.
Now we recall some important results of [TW] and [Mu] that relate to our work.
Theorem 2.1 ([TW, Theorem 1.4] [Mu, Theorem 3.2])
Let $(L,\leq)$ be a finite semidistributive lattice. Then the following are equivalent:
@@
We note that, without semidistributvity, the notions of extremality and left modularity do not imply each other. More explicitly, there exist extremal lattices that are not left modular, and there also exist left modular lattices that are not extremal (see [TW, Figure 5]). Moreover, as shown in [TW, Figure 3], there exist trim lattices that are not semidistributive, and there also exist semidistributive lattices that are not trim.
@@ 2.2. Torsion theories
    $\mathcal{T}$ and $\mathcal{F}$ have only 0 in common.
@@ 2.4. Chain of bricks
We note that the map $f$ in the above theorem is almost never invertible. In fact, $f$ is bijective if and only if $A$ is local. We remark that starting from a chain of bricks $\{B_{i}\}_{i\in I}$ in $\operatorname*{ch-brick}A$, we often have that the cardinality of the corresponding chain of torsion classes in $\operatorname*{ch-tors}A$ induced by $f$ is no smaller (and often larger) than the cardinality of $I$. Moreover, one immediately concludes that the length of each chain in $\operatorname*{tors}A$ is bounded by $|{\operatorname*{brick}A}|$. Also, note that $A$ is brick-infinite if and only if $\operatorname*{tors}A$ contains an infinite chain. For more details on the chain of torsion classes and their interactions with bricks, see [BCZ], [DIJ], [DI+], and [KD, Appendix].
@@ 3. Brick-splitting torsion pairs
3.1. Characterizations
Proposition 3.1
  1. (1)
    $(\mathcal{T},\mathcal{T}^{\perp})$ is a brick-splitting torsion pair;
  2. (2)
Proof.
Assume that (1) holds. Consider $X\in\operatorname*{brick}A$. Since $(\mathcal{T},\mathcal{T}^{\perp})$ is brick-splitting, we get $X\in\mathcal{T}$ or $X\in\mathcal{T}^{\perp}$. In the first case, $\mathcal{T}(X)$ is a torsion class in the interval $0,\mathcal{T}]$, and we have $X$ is the labeling brick of the minimal inclusion $\mathcal{T}(X)\cap{}^{\perp}X\subset\mathcal{T}(X)$. In the second case, if $X\in\mathcal{T}^{\perp}$, then $X$ is the labeling brick of the minimal inclusion $\mathcal{F}(X)\cap X^{\perp}\subset\mathcal{F}(X)$ of torsion-free classes, all included in $\mathcal{T}^{\perp}$. This translates to $X$ being a brick-labeling of the corresponding inclusion $^{\perp}X\subset{{}^{\perp}(\mathcal{F}(X)\cap X^{\perp}})$ of torsion classes in $\mathcal{T},\operatorname*{mod}A]$.
Now, assume that $(2)$ holds. Let $X\in\operatorname*{brick}A$. If $X$ is a labeling brick between two torsion classes in $0,\mathcal{T}]$, then obviously $X\in\mathcal{T}$. Suppose that $X$ is the labeling brick between some two adjacent torsion classes $\mathcal{T}_{1}\subset\mathcal{T}_{2}$ in the interval $\mathcal{T},\operatorname*{mod}A]$. This implies that $X\in\mathcal{T}_{1}^{\perp}\subseteq\mathcal{T}^{\perp}$. ∎
Remark 3.2
For brick-finite algebras, the above proposition gives an easy criterion to verify whether a given torsion class admits a brick-splitting torsion pair. For instance, in Example 3.12, one can check that every $X\in\operatorname*{brick}A$ appears on the left chain of the lattice $\operatorname*{tors}A$ (see Figure 1). Consequently, every torsion class on this chain gives rise to a brick-splitting torsion pair. In contrast, the labeling bricks on the other maximal chain on the right side consists of only two bricks, whereas $\operatorname*{brick}A$ has three elements. Thus, we can immediately conclude that $\mathcal{T}^{\prime\prime}=\operatorname{Fac}(\begin{smallmatrix}2\\ 2\end{smallmatrix})$ is not a brick-splitting torsion class. In particular, ${\begin{smallmatrix}1\\ 2\end{smallmatrix}}\in\operatorname*{brick}A$ does not appear in $0,\mathcal{T}^{\prime\prime}]$ nor $\mathcal{T}^{\prime\prime},\operatorname*{mod}A]$.
Corollary 3.3
Proof.
Proposition 3.4
  1. (1)
    $(\mathcal{T},\mathcal{T}^{\perp})$ is a brick-splitting torsion pair;
  2. (2)
Proof.
$0\to T\to L\to F\to 0$
for $L$ where $T\in\mathcal{T}$ and $F\in\mathcal{T}^{\perp}$. Since $L$ is a minimal extending brick for $\mathcal{Y}$, we know that every proper quotient of $L$ has to be in $\mathcal{Y}$. So if $T$ is non-zero, then $F\in\mathcal{Y}$, making $L\in\mathcal{Y}\vee\mathcal{T}$, a contradiction. Therefore, $T=0$ and $L\in\mathcal{T}^{\perp}$.
For the other direction, assume that $\mathcal{T}$ is brick-splitting. Observe that it follows from [DI+, Theorem 1.3] that $\operatorname*{tors}A$ is weakly atomic. By the proof of [TW, Lemma 2.8] and weakly atomic property of $\operatorname*{tors}A$, in order to prove that $\mathcal{T}$ is left modular, it suffices to prove that for two torsion classes $\mathcal{V}\subsetneq\mathcal{U}$ forming a cover relation, we have
Observe that the containment from left to right is clear. Now, let $L$ be the labeling brick corresponding to the cover relation $\mathcal{V}\subsetneq\mathcal{U}$. Then we have $\mathcal{T}(L)\vee\mathcal{V}=\mathcal{U}$ and $^{\perp}L\cap\mathcal{U}=\mathcal{V}$. Since $\mathcal{T}$ is brick-splitting, either $L\in\mathcal{T}$ or $L\in\mathcal{T}^{\perp}$. If $L\in\mathcal{T}$, we see that
$(\mathcal{U}\cap\mathcal{T})\vee\mathcal{V}\supseteq\mathcal{T}(L)\vee\mathcal{V}=\mathcal{U}\supseteq\mathcal{U}\cap(\mathcal{T}\vee\mathcal{V}).$
Assume now that $L\in\mathcal{T}^{\perp}$. Since both $\mathcal{T}$ and $\mathcal{V}$ are contained in $^{\perp}L$, we have
$\mathcal{U}\cap(\mathcal{T}\vee\mathcal{V})\subseteq\mathcal{U}\cap{{}^{\perp}L}=\mathcal{V}\subseteq(\mathcal{U}\cap\mathcal{T})\vee\mathcal{V}.$
Therefore, ()(*) also holds in this case. ∎
3.2. Brick quivers and brick-splitting torsion classes
To any algebra, one can explicitly associate a quiver that captures the interactions between the bricks and hence the brick-directedness of the algebra under consideration. In particular, for a connected algebra $A$, by $Q^{b}(A)$ we denote the brick-quiver of $A$, constructed as follows: The vertices of $Q^{b}(A)$ are in bijection with elements of $\operatorname*{brick}A$, and for each pair of distinct (hence non-isomorphic) modules $X$ and $Y$ in $\operatorname*{brick}A$, we put an arrow from $X$ to $Y$ if $\operatorname{Hom}_{A}(X,Y)\neq 0$. For an explicit example, see Example 4.11. One motivation for studying the brick-quiver is the following.
Proposition 3.5
There exists a bijection between the following sets.
  1. (1)
    Brick-splitting torsion classes of $A$;
  2. (2)
    Successor-closed subsets of $Q^{b}(A)$.
Proof.
Given a brick-splitting torsion class $\mathcal{T}$, we associate to it the set $\mathcal{T}\cap\operatorname*{brick}A$, which is regarded as a set of vertices of $Q^{b}(A)$. If $B\to B^{\prime}$ is an arrow in $Q^{b}(A)$ with $B\in\mathcal{T}$, then $\operatorname{Hom}_{A}(B,B^{\prime})\neq 0$ so that $B^{\prime}\not\in\mathcal{T}^{\perp}$. Since $\mathcal{T}$ is brick-splitting, this yields that $B^{\prime}\in\mathcal{T}$. Hence, the set $\operatorname*{brick}(A)\cap\mathcal{T}$ is successor-closed in $Q^{b}(A)$. Since torsion classes are determined by the bricks they contain, the assignment $\mathcal{T}\mapsto\mathcal{T}\cap\operatorname*{brick}A$ is injective.
Conversely, let $\mathcal{S}$ be a set of bricks that forms a successor-closed set in $Q^{b}(A)$. We let $\mathcal{T}=\mathcal{T}(\mathcal{S})$ be the smallest torsion class that contains $S$. We observe that for a brick $B$, we have $\operatorname{Hom}_{A}(\mathcal{S},B)=0$ if and only if $B\in\mathcal{T}^{\perp}$, if and only if $B\not\in\mathcal{S}$, where the latter equivalence follows from the fact that $\mathcal{S}$ is successor-closed in $Q^{b}(A)$. It follows that $\mathcal{T}\cap\operatorname*{brick}A=\mathcal{S}$. Finally, we check that the constructed $\mathcal{T}$ is brick-splitting. Let $B\in\operatorname*{brick}A$. If $B\in\mathcal{S}$, then $B\in\mathcal{T}$. Otherwise, it follows from what we have shown that $B\in\mathcal{T}^{\perp}$. Hence, $\mathcal{T}$ is brick-splitting. ∎
Proposition 3.6
  1. (1)
    $Q^{b}(A)$ is connected and has no loops.
  2. (2)
  3. (3)
  4. (4)
  5. (5)
Proof.
  1. (1)
    $0\to S_{j}\to M\to S_{i}\to 0.$
  2. (2)
  3. (3)
  4. (4)
    This directly follows from the definition.
  5. (5)
Remark 3.7
As observed in [MP2], from some earlier results of [As1], it follows that for any brick-finite algebra $A$ of rank $n$, there is no set of pairwise Hom-orthogonal modules of size $n+1$. In particular, for such an algebra, the maximal size of an independent set in $Q^{b}(A)$ is $n$, that is, the independence number of $Q^{b}(A)$ is $n$ and is attained exactly for the set of vertices in $Q^{b}(A)$ corresponding to the simple $A$-modules (see [MP2, Theorem A]). In light of the Semibrick conjecture (see [MP2, Conjecture 2.2]), it is expected that the converse of the latter statement is true, that is, the independence number of $Q^{b}(A)$ is $n$ if and only if $A$ is brick-finite. This is already known to be the case for various families of algebras, but it remains open in full generality (for more details, see [MP2, MP4] and references therein).